OpenStudy (anonymous):

Pam likes to practice dancing while preparing for a math tournament. She spends 80 minutes every day practicing dance and math. To help her concentrate better on math, she spends 20 minutes more dancing than doing math. Part A: Write a pair of linear equations to show the relationship between the number of minutes Pam practices math every day (x) and the number of minutes she practices dances every day (y). (5 points) Part B: How much time does Pam spend on practicing math every day? (3 points) Part C: Is it possible for Pam to have spent 60 minutes practicing dance every day? Expla

OpenStudy (anonymous):

^^Explain your reasoning

OpenStudy (isaidavila):

Part A) I think could be x+y=80 y=x+20

OpenStudy (isaidavila):

Solving this system of equations you can figure out the number of minutes she spends on math and dancing

OpenStudy (anonymous):

r u sure on part a bc i cant get it wrong

OpenStudy (anonymous):

@Kainui

OpenStudy (isaidavila):

The first equation is right because the number of math and dancing = 80

OpenStudy (isaidavila):

Better have another opinion

OpenStudy (anonymous):

ok do u know someone we can tag?

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@ILUVSOCCER

OpenStudy (anonymous):

can u help plz!! @ILUVSOCCER

OpenStudy (iluvsoccer):

I'll try.

OpenStudy (mertsj):

Solve the system you have been given by the previous helper and see if the answers work in the problem description.

OpenStudy (anonymous):

ive tried but could u help ?

OpenStudy (anonymous):

is this right so far?

OpenStudy (anonymous):

Part A: x+y=80 y=x+20 Part B: Pam spends 20 practicing math everyday The reason why is because if we divided 80/2 it would be 40 but add 20 on dance and subtract 20 on math so dance would have 60 minutes and math would have 20 minutes.

OpenStudy (anonymous):

@ILUVSOCCER and @Mertsj

OpenStudy (mertsj):

Replace y in this equation: x+y=80 with y = x+20

OpenStudy (mertsj):

Now you have: x + x+20 =80 Solve that equation.

OpenStudy (anonymous):

thts not what im supposed to do i need x and y

OpenStudy (isaidavila):

Solve for x and replace angain in any equation to have the value of y

OpenStudy (mertsj):

You have the equation x + x+20=80. If you solve that equation you will have x

OpenStudy (anonymous):

i finished tht one could u help on another?

OpenStudy (anonymous):

.Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B: Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between −3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points

OpenStudy (isaidavila):

Part A: To find the intersection points of two equation you have to match both equations. So that why 8x=2x+2 is the solution for 8x=2x+2

OpenStudy (anonymous):

thx do u know part b and c?

OpenStudy (anonymous):

@isaidavila

OpenStudy (isaidavila):

I'm not so sure in part be because if you solve for x you got 1/3

OpenStudy (anonymous):

@ILUVSOCCER would u know?

OpenStudy (anonymous):

@Study312 @raffle_snaffle @jhonyy9

OpenStudy (study312):

What is the question? I mean she spends 60 mins doing math because is 80 is the total of dance and math and 20 is of dance then math is 60

OpenStudy (anonymous):

no i wrote the 2nd one is what we are on

OpenStudy (anonymous):

do u see it @study312

OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2x + 2 intersect are the solutions of the equation 8x = 2x + 2. (4 points) Part B: Make tables to find the solution to 8x = 2x + 2. Take the integer values of x between −3 and 3. (4 points) Part C: How can you solve the equation 8x = 2x + 2 graphically? (2 points)

OpenStudy (anonymous):

@jhonyy9

OpenStudy (study312):

idk how to do that one

OpenStudy (anonymous):

how about this one?

OpenStudy (anonymous):

Graph the system of inequalities presented here on your own paper, then use your graph to answer the following questions: y > 2x + 3 y is less than negative 3 over 2 times x minus 4 Part A: Describe the graph of the system, including shading and the types of lines graphed. Provide a description of the solution area. (6 points) Part B: Is the point (−4, 6) included in the solution area for the system? Justify your answer mathematically. (4 points)

OpenStudy (anonymous):

@Study312

OpenStudy (study312):

sorry not off the top of my head

OpenStudy (anonymous):

@Ms-Brains

OpenStudy (anonymous):

@isaidavila

OpenStudy (anonymous):

@Data_LG2 WILL MEDAL AND FAN!!!